HUBBLE SPACE TELESCOPE DAILY REPORT #3002 PERIOD COVERED: 0000Z (UTC) 11/20/01 - 0000Z (UTC) 11/21/01 Daily Status Report as of 325/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 9128 (A Snapshot Survey of Absorption Lines from High Velocity Clouds in the Milky Way) The Space Telescope Imaging Spectrograph (CCD) was used to search for Mg II and Mg I absorption from Galactic High Velocity Clouds {HVCs} along the sightlines of 51 of the brightest QSOs, BL Lacs and Seyfert 2 galaxies in the sky. The observations completed with no reported problems. 1.2 Completed Nine Sets of WF/PC-2 8942 (Cycle 10 Intflat Sweeps and Linearity Test) The WF/PC-2 was used to monitor the pixel-to-pixel flatfield response and provide a linearity check. These intflat sequences are done once during the year and the images will provide a backup database in the event of complete failure of the visflat lamp as well as allow monitoring of the gain ratios. There were no reported problems. 1.3 Completed Two Sets of STIS/CCD 8901 (Dark Monitor-Part 1) ! The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 1.4 Completed STIS/CCD 8611 (Observations of Nearby Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to obtain UV spectra of five nearby {0.02 < z < 0.08} SNe Ia in the Hubble Flow. The spectra will be taken at weekly intervals over a range in time starting slightly before maximum light and extending to +30 days. These observations will accomplish the following three goals: {1} calibration of the rest frame UV light curves of SNe Ia and an assessment of their potential use as distance indicators through UV light curve shape analyses. {2} improvement in our understanding of the physics of SNe Ia, metallicity/evolutionary effects and correlations between peak brightness and UV spectral features. {3} calibration of the SNe Ia previously observed by HST at high-redshift. This data is crucial for proper cross-filter k-corrections and calibration of the supernova photometry. The observations completed with no reported anomalous activity. 1.5 Completed Three Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.6 Completed STIS/CCD/MA2 8569 (A New Survey for Low-Redshift Damped Lyman-Alpha Lines in QSO MgII Systems) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to support studies which have shown that most of the observable neutral gas mass in the Universe resides in QSO damped LyAlpha {DLA} systems. However, at low redshift, DLA can only be found by searching in the UV with HST. By boot-strapping from the MgII statistics, we will be able to further improve the determination of the low- redshift statistical properties of DLA {their incidence and cosmological mass density} and open up new opportunities for studies at low redshift. The observations completed nominally. 1.7 Completed STIS/CCD 8903 (Bias Monitor - Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 to build up high-S/N superbiases and track the evolution of hot columns. There were no problems. 1.8 Completed WF/PC-2/STIS/CCD 8324 (The Structure and Kinematics of Irradiated Disks and Associated High Velocity Features in Orion) Both the WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to make observations of high velocity feature M-42-LV2-LV located within the Orion Nebula. The STIS provides us with a unique opportunity to study the early evolution of proto-planetary disks and their chances of survival in the harsh irradiated environments of OB class associations where most stars form. A wavelength calibration observation was also taken. The observations were completed as planned, and no anomalies were reported. 1.9 Completed WF/PC-2 9118 (Tracing the Cosmic Expansion to z>1 with Type Ia Supernovae) The WF/PC-2 was used to observe type Ia supernovae in order to provide evidence for an accelerating universe. The case for cosmic acceleration rests almost entirely on the observation that the observed SN Ia at z~0.5 are 0.25 magnitudes fainter than expected for a non-accelerating Universe. It is proposed to follow five SN Ia in the range 0.95 0 cosmology, this experiment is a powerful and straightforward way to assess the reliability of the SN Ia measurements. In addition, if SN Ia are reliable standard candles, the proposed observations will significantly increase the precision with which Omega_Lambda and Omega_M are measured. There was no reported anomalous activity. 1.10 Completed Nine Sets of STIS/CCD 8910 (Sparse Field CTE Test-Internal {Cycle 10}) The Space Telescope Imaging Spectrograph (CCD) was used to perform CTE measurements that are made using the "sparse field test", along both the serial and parallel axes. This program needs special commanding to provide {a} off-center MSM positionings of some slits, and {b} the ability to read out with any amplifier {A, B, C, or D}. No problems were encountered. 1.11 Completed Three Sets of WF/PC-2 8936 (Cycle 10 Supplemental Darks Pt1/3) The WF/PC-2 was used to perform a dark calibration program that obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 6 Successful: 6 Scheduled Re-acquisitions: 10 Successful: 10 2.2 FHST Updates: Scheduled: 13 Successful: 13 2.3 Operations Notes: The SSR EDAC error counter was cleared at 325/0133Z per ROP SR-1A. A TTR was generated when a re-transmit was required at 325/024215Z during a NSSC-1 load. Accordingly, at 325/0244Z, using ROP NS-5, SI C&DH errors were reset. The SSR-3 pointer was set at 325/0320Z as directed by ROP SR-3. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW